### v1

# InFL-UX: A Toolkit for Web-Based Interactive Federated Learning

Abstract.
This paper presents InFL-UX, an interactive, proof-of-concept browser-based Federated Learning (FL) toolkit designed to integrate user contributions seamlessly into the machine learning (ML) workflow. InFL-UX enables users across multiple devices to upload datasets, define classes, and collaboratively train classification models directly in the browser using modern web technologies. Unlike traditional FL toolkits, which often focus on backend simulations, InFL-UX provides a simple user interface for researchers to explore how users interact with and contribute to FL systems in real-world, interactive settings. By prioritising usability and decentralised model training, InFL-UX bridges the gap between FL and Interactive Machine Learning (IML), empowering non-technical users to actively participate in ML classification tasks.
1. Introduction and Related Work
Traditional machine learning (ML) often struggles with limited data, particularly in specialised domains where data collection is costly or labour-intensive (Roh et al., 2021). The lack of direct involvement of domain experts, due to technical barriers, further delays the acquisition of new training data (Holstein, 2024). To address these challenges, Fails and Olsen (2003) introduced interactive machine learning (IML), enabling non-technical users to train ML models using their own data through manual classification or correcting model outputs. Unlike traditional ML, IML allows real-time updates in response to user input, facilitating focused and incremental adjustments (Amershi et al., 2014; Dudley and Kristensson, 2018). Building on these advancements, Tseng et al. (2023) developed Co-ML, a tablet-based application for collaboratively building ML image classification models across multiple devices, focusing on teaching dataset design practices by creating a shared dataset. In this paper, we extend these concepts by proposing a browser-based tool that allows users to collaborate on IML tasks using federated learning.
Typically, ML models are trained on centralised datasets. However, in domains such as healthcare, data is often distributed across multiple devices and cannot be shared due to privacy concerns. Federated learning (FL) addresses this limitation by enabling decentralised training of a shared model while keeping data on the clients’ devices (McMahan et al., 2017). FL operates by aggregating locally computed updates to a central server. Two key types of FL algorithms exist: synchronous and asynchronous. Synchronous FL (McMahan et al., 2017) requires all participating clients to train simultaneously after receiving the latest global model, submitting updates together for incorporation into the global model. In contrast, asynchronous FL (Xie et al., 2019) allows clients to train and submit updates independently, with the server updating the global model upon each arrival. Given its flexibility, asynchronous FL is more suitable for our use case. Recent advancements in JavaScript-based deep learning (DL) frameworks, such as TensorFlow.js (Smilkov et al., 2019), have made browser-based DL feasible (Ma et al., 2019). Google’s Teachable Machine (Carney et al., 2020) demonstrates this with a no-code web application that enables users to train ML models locally via an intuitive interface. However, this tool focuses exclusively on local training without integrating FL capabilities. In contrast, browser-based FL frameworks, such as those by Lian et al. (2022) and Ángel Morell and Alba (2022), support FL but lack interactivity, requiring users to provide data in predefined folders with minimal engagement.
In this paper, we introduce InFL-UX111The name is a combination of Interactive Machine Learning, Federated Learning, and User Experience., an interactive, browser-based FL toolkit that demonstrates a proof-of-concept (POC) implementation leveraging modern web technologies. The application employs asynchronous FL, implemented using FedAsync (Xie et al., 2019), a method chosen for its simplicity and single-hyperparameter configuration, enabling collaborative model training. Users can upload datasets, define custom classes, and train ML models tailored to diverse classification tasks. While most FL research focuses on simulations, InFL-UX combines FL and IML, prioritising user engagement to address FL challenges from a user-centric perspective. Our goal is to assist FL practitioners in integrating FL setups into intelligent user interfaces and evaluating them in user studies. Future development will enhance the system with advanced FL aggregation strategies and ML tasks, such as data annotation, to increase adoption across various domains.
2. System Design
InFL-UX utilises ONNX Runtime222https://onnxruntime.ai/ (Accessed January 03, 2025), a cross-platform machine learning library designed for fast on-device inference and training, with support for web browsers. It leverages modern browser APIs such as WebAssembly (Rossberg, 2022), a low-level, assembly-like language enabling near-native performance for web applications, and WebGPU (Ninomiya et al., 2024), which allows high-performance GPU computations and improves compatibility with modern GPUs as the successor to WebGL333https://www.khronos.org/webgl/ (Accessed January 03, 2025). This browser-based implementation facilitates easy integration with existing systems due to the ubiquity of web browsers and offers seamless compatibility with models from popular ML frameworks through the ONNX (Open Neural Network Exchange) format. The application is developed in Python using Flask444https://flask.palletsprojects.com/en/stable/ (Accessed January 03, 2025), and integrates ONNX Runtime via JavaScript and encapsulates UI components using Web Components555https://www.webcomponents.org/ (Accessed January 03, 2025), ensuring independence from specific UI frameworks. Persistent browser storage is achieved through IndexedDB (Bell, 2024), enabling the retention of user data, such as uploaded files and inference results, between sessions, even after the browser is closed. The code for InFL-UX is publicly available online: https://github.com/tmaurer42/interactive-fl-poc.
2.1. System and Application Design
The system architecture, depicted in Figure 1, consists of two main components: the FL Server and the FL Web Server. The FL Web Server hosts the web application and delivers necessary files to client devices, while the FL Server manages client updates, maintains the global model, and provides the latest model version and training instructions upon request. This architecture underpins the POC application, which focuses on image classification but can also support tasks like object detection and image segmentation via abstract classes. The system enables the simulation of multiple independent clients to demonstrate its functionality. One client acts as the administrator, configuring the setup through the configuration page (Figure 2a) by defining the use case, selecting the aggregator and ML model, and setting training parameters. On the client side, users upload images, receive suggested labels based on the current model, and can accept or modify these labels (Figure 2b). After reviewing, a training session can be initiated using the local dataset (Figure 2c). Upon training completion, updates are sent automatically to the central server for aggregation. Clients can also test the global model using their own testing datasets (Figure 2d).
Currently, the system supports a single FL aggregation strategy, FedAsync, and one ML model, MobileNetV2 (Sandler et al., 2019). Extending the system to include a new aggregation strategy or ML model is straightforward and requires minimal code changes. A new aggregation strategy involves defining the strategy as a Python function and adding its configurable parameters to the administrator’s configuration interface. Similarly, new ML models can be incorporated by extending the administrator interface’s dropdown menu with additional PyTorch models666https://pytorch.org/ (Accessed January 03, 2025). The system is deployed using Docker777https://www.docker.com/ (Accessed January 03, 2025). InFL-UX is designed to be extensible, enabling the integration of additional aggregation strategies. However, it does not support simulation tasks, as these are handled by established frameworks such as Flower (Beutel et al., 2022). Instead, InFL-UX serves as a subsequent step following the simulation phase, focusing on validating aggregation methods in a real-world context.
2.2. Limitations
Several limitations emerged during the development of the application. The ONNX Runtime integration for the web, being relatively new, has incomplete functionalities for model training. For instance, the learning rate of the optimiser cannot be adjusted, requiring clients to use a default rate of 0.001, and only a limited set of loss functions are available. Additionally, WebGPU currently supports only inference, while training is restricted to the WebAssembly backend. Browser-specific constraints also affect the application. The size of the local dataset stored in the browser’s IndexedDB is restricted, with storage limits varying by browser and disk space availability. Furthermore, the WebGPU API is currently accessible only in developer builds of modern browsers, which are often disallowed on systems with strict security policies.
3. Conclusion and Future Work
The development of the POC InFL-UX demonstrated the feasibility of implementing an interactive, browser-based FL system, integrating state-of-the-art technologies from the UI to the server-side aggregation process. The approach opens the door for integrating FL capabilities into IML systems, especially in privacy-sensitive domains. Despite its limitations, the application successfully utilised ONNX Runtime and modern browser features, such as WebAssembly and WebGPU, to enable client-side training for FL. The framework supports collaborative training across devices while maintaining compatibility and promoting user adoption.
Future work should focus on conducting user studies to evaluate adoption and performance in real-world scenarios. Expanding the application to include additional computer vision tasks, such as image segmentation or object detection, is another promising direction. Integrating explainable AI techniques, such as class activation maps (Selvaraju et al., 2017), could enhance transparency and user trust. Additionally, enabling administrators to deploy new training tasks and allowing clients to participate in multiple tasks with distinct datasets for each use case would significantly broaden the application’s utility.
Acknowledgements.
This work was funded, in part, by the German Federal Ministry of Education and Research (BMBF) under grant number 01IW23002 (No-IDLE).References
- (1)
- Amershi et al. (2014) Saleema Amershi, Maya Cakmak, William Bradley Knox, and Todd Kulesza. 2014. Power to the People: The Role of Humans in Interactive Machine Learning. AI Magazine 35, 4 (Dec. 2014), 105–120. https://doi.org/10.1609/aimag.v35i4.2513
- Bell (2024) Joshua Bell. 2024. Indexed Database API 3.0. Technical Report. W3C. https://www.w3.org/TR/IndexedDB/ Version Number: 3.0.
- Beutel et al. (2022) Daniel J. Beutel, Taner Topal, Akhil Mathur, Xinchi Qiu, Javier Fernandez-Marques, Yan Gao, Lorenzo Sani, Kwing Hei Li, Titouan Parcollet, Pedro Porto Buarque de Gusmão, and Nicholas D. Lane. 2022. Flower: A Friendly Federated Learning Research Framework. https://doi.org/10.48550/arXiv.2007.14390 arXiv:2007.14390 [cs].
- Carney et al. (2020) Michelle Carney, Barron Webster, Irene Alvarado, Kyle Phillips, Noura Howell, Jordan Griffith, Jonas Jongejan, Amit Pitaru, and Alexander Chen. 2020. Teachable Machine: Approachable Web-Based Tool for Exploring Machine Learning Classification. In Extended Abstracts of the 2020 CHI Conference on Human Factors in Computing Systems (¡conf-loc¿, ¡city¿Honolulu¡/city¿, ¡state¿HI¡/state¿, ¡country¿USA¡/country¿, ¡/conf-loc¿) (CHI EA ’20). Association for Computing Machinery, New York, NY, USA, 1–8. https://doi.org/10.1145/3334480.3382839
- Dudley and Kristensson (2018) John J. Dudley and Per Ola Kristensson. 2018. A Review of User Interface Design for Interactive Machine Learning. ACM Transactions on Interactive Intelligent Systems 8, 2 (June 2018), 1–37. https://doi.org/10.1145/3185517
- Fails and Olsen (2003) Jerry Alan Fails and Dan R. Olsen. 2003. Interactive machine learning. In Proceedings of the 8th International Conference on Intelligent User Interfaces (Miami, Florida, USA) (IUI ’03). Association for Computing Machinery, New York, NY, USA, 39–45. https://doi.org/10.1145/604045.604056
- Holstein (2024) Joshua Holstein. 2024. Bridging Domain Expertise and AI through Data Understanding. In Companion Proceedings of the 29th International Conference on Intelligent User Interfaces (IUI ’24 Companion). Association for Computing Machinery, New York, NY, USA, 163–165. https://doi.org/10.1145/3640544.3645248
- Lian et al. (2022) Zhuotao Lian, Qinglin Yang, Qingkui Zeng, and Chunhua Su. 2022. WebFed: Cross-platform Federated Learning Framework Based on Web Browser with Local Differential Privacy. In ICC 2022 - IEEE International Conference on Communications. 2071–2076. https://doi.org/10.1109/ICC45855.2022.9838421
- Ma et al. (2019) Yun Ma, Dongwei Xiang, Shuyu Zheng, Deyu Tian, and Xuanzhe Liu. 2019. Moving Deep Learning into Web Browser: How Far Can We Go?. In The World Wide Web Conference (San Francisco, CA, USA) (WWW ’19). Association for Computing Machinery, New York, NY, USA, 1234–1244. https://doi.org/10.1145/3308558.3313639
- McMahan et al. (2017) Brendan McMahan, Eider Moore, Daniel Ramage, Seth Hampson, and Blaise Aguera y Arcas. 2017. Communication-Efficient Learning of Deep Networks from Decentralized Data. In Proceedings of the 20th International Conference on Artificial Intelligence and Statistics (Proceedings of Machine Learning Research, Vol. 54), Aarti Singh and Jerry Zhu (Eds.). PMLR, 1273–1282. https://doi.org/10.48550/arXiv.1602.05629
- Ninomiya et al. (2024) Kai Ninomiya, Jim Blandy, and Brandon Jones. 2024. WebGPU. W3C Working Draft. W3C. https://www.w3.org/TR/webgpu/
- Roh et al. (2021) Yuji Roh, Geon Heo, and Steven Euijong Whang. 2021. A Survey on Data Collection for Machine Learning: A Big Data - AI Integration Perspective. IEEE Transactions on Knowledge and Data Engineering 33, 4 (April 2021), 1328–1347. https://doi.org/10.1109/TKDE.2019.2946162 Conference Name: IEEE Transactions on Knowledge and Data Engineering.
- Rossberg (2022) Andreas Rossberg. 2022. WebAssembly Core Specification. Technical Report. W3C. https://www.w3.org/TR/wasm-core-2/ Version Number: 2.0.
- Sandler et al. (2019) Mark Sandler, Andrew Howard, Menglong Zhu, Andrey Zhmoginov, and Liang-Chieh Chen. 2019. MobileNetV2: Inverted Residuals and Linear Bottlenecks. https://doi.org/10.48550/arXiv.1801.04381 arXiv:1801.04381 [cs].
- Selvaraju et al. (2017) Ramprasaath R. Selvaraju, Michael Cogswell, Abhishek Das, Ramakrishna Vedantam, Devi Parikh, and Dhruv Batra. 2017. Grad-CAM: Visual Explanations from Deep Networks via Gradient-Based Localization. In 2017 IEEE International Conference on Computer Vision (ICCV). 618–626. https://doi.org/10.1109/ICCV.2017.74 ISSN: 2380-7504.
- Smilkov et al. (2019) Daniel Smilkov, Nikhil Thorat, Yannick Assogba, Charles Nicholson, Nick Kreeger, Ping Yu, Shanqing Cai, Eric Nielsen, David Soegel, Stan Bileschi, Michael Terry, Ann Yuan, Kangyi Zhang, Sandeep Gupta, Sarah Sirajuddin, D Sculley, Rajat Monga, Greg Corrado, Fernanda Viegas, and Martin M Wattenberg. 2019. TensorFlow.js: Machine Learning For The Web and Beyond. In Proceedings of Machine Learning and Systems, A. Talwalkar, V. Smith, and M. Zaharia (Eds.), Vol. 1. 309–321. https://doi.org/10.48550/arXiv.1901.05350
- Tseng et al. (2023) Tiffany Tseng, Jennifer King Chen, Mona Abdelrahman, Mary Beth Kery, Fred Hohman, Adriana Hilliard, and R. Benjamin Shapiro. 2023. Collaborative Machine Learning Model Building with Families Using Co-ML. In Proceedings of the 22nd Annual ACM Interaction Design and Children Conference (Chicago, IL, USA) (IDC ’23). Association for Computing Machinery, New York, NY, USA, 40–51. https://doi.org/10.1145/3585088.3589356
- Xie et al. (2019) Cong Xie, Sanmi Koyejo, and Indranil Gupta. 2019. Asynchronous Federated Optimization. arXiv e-prints, Article arXiv:1903.03934 (March 2019). https://doi.org/10.48550/arXiv.1903.03934 arXiv:1903.03934 [cs.DC]
- Ángel Morell and Alba (2022) José Ángel Morell and Enrique Alba. 2022. Dynamic and adaptive fault-tolerant asynchronous federated learning using volunteer edge devices. Future Generation Computer Systems 133 (2022), 53–67. https://doi.org/10.1016/j.future.2022.02.024